OpenStudy (anonymous):

What is the coefficient of rolling friction for the tire under low pressure?

OpenStudy (anonymous):

Two bicycle tires are set rolling with the same initial speed of 4.00 along a long, straight road, and the distance each travels before its speed is reduced by half is measured. One tire is inflated to a pressure of 40 psi and goes 18.7 ; the other is at 105 psi and goes 93.1 . Assume that the net horizontal force is due to rolling friction only.

OpenStudy (anonymous):

find 'a' for the first tire from: (Vi/2)^2 = Vi^2 -2ad -(3/4)Vi^2 = -2ad Ff = ma = mu*mg a=mu*g mu = a/g